OpenStudy (datanewb):

Question regarding a summation equality: \[\sum_{k=0}^n n(n-1)\dotsb(n-k+1) \leq \sum_{k=0}^n n^k \]

OpenStudy (datanewb):

I don't necessarily need a rigorous proof (although I wouldn't mind that), but how can I show myself that the above equality is true for all $n$? Well, I'm seeing it now. on the left and right side of the equation we will have on average n+k2 terms for each addend. On the right instead of counting down from n to n-k+1, each of those terms is equal to n. On the left each term is less than or equal to n... Makes sense now.

OpenStudy (anonymous):

\[\sum_{k=0}^{n}n^2-kn+n \le \sum_{k=0}^{n}n^k\] take log of both sides \[\sum_{k=0}^{n}2\ln n-\ln(kn)+\ln n \le \sum_{k=0}^{n}k \ln n\] factor out ln n \[\sum_{k=0}^{n}\ln n \left( 2-1+\ln k+1 \right)\le \sum_{k=0}^{n}k \ln n\] Divide by ln n \[\sum_{k=0}^{n} 2+\ln k \le \sum_{k=0}^{n}k\] Get rid of the summation notation \[2+\ln k \le k\] and just solve the inequality

OpenStudy (datanewb):

That is very good, but how do I show that \[ \sum_{k=0}^n n(n-1)\dotsb(n-k+1) \leq \sum_{k=0}^{n}n^2-kn+n \] Also, I think the step after "factor out ln n" should be: \[\sum_{k=0}^{n}\ln n \left( 2-1-\ln k+1 \right)\le \sum_{k=0}^{n}k \ln n\] which would give \[2-\ln k \le k\] so by that logic, the summation is true for integer values of n > 1, although I don't quite follow the very first step.